OpenStudy (anonymous):

Can anyone help me on this Geometry assignment? It is my worst subject.... There are 2 assignments that I have to turn in by Sunday. My brain is fried. I don't know what else to do. HELP!

OpenStudy (mathmale):

Please post the assignment and let us know what kind of help you need. Welcome to OpenStudy!!

OpenStudy (anonymous):

Trigonometric Functions 1.Draw a right triangle. Roll a die two times and record the numbers you roll. On the drawing of your right triangle, label the legs of the triangle using the numbers you rolled. Use the Pythagorean Theorem to find the length of the hypotenuse and write out the steps you performed using words. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: If a seven and nine are rolled, then the legs of the triangle can be labeled as follows: 2.Draw another right triangle. Roll a die one time and record the number you roll. Label one of the legs of a right triangle with the number you rolled. Now, roll the die two more times and record the numbers you roll. Use the two numbers you rolled to create an angle measure for your triangle. On the drawing of your right triangle, label the leg and angle of the triangle using the numbers you rolled. Find the length of the missing leg and find the missing angle measure. Write a paragraph telling how you found the missing values. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: Roll # 1: A die is shown with the number 11 visible. Rolls # 2 and # 3: A picture of two dice is shown. One die is showing the number 7 and the other die is showing the number 9. If an eleven is rolled, then label one leg of the triangle eleven units in length. If the next two rolls yield a seven and nine, then label one angle as 79°. A right triangle is shown. One leg is labeled 11 and an interior angle is labeled 79 degrees. 3.Draw another right triangle. Roll a die two times and record the numbers you roll. On the drawing of your right triangle, label the legs of the triangle using the numbers you rolled. Use the Pythagorean Theorem and trigonometric functions to find the length of the hypotenuse and the missing angle measures. Write a paragraph telling how you found the missing values. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: Rolls # 1 and # 2: Two dice are shown. One die is showing the number 9 and the other die is showing the number 11.

OpenStudy (anonymous):

Its a load of gibberish to me, can you help walk me through it?

OpenStudy (anonymous):

1.Draw a right triangle. Roll a die two times and record the numbers you roll. On the drawing of your right triangle, label the legs of the triangle using the numbers you rolled. Use the Pythagorean Theorem to find the length of the hypotenuse and write out the steps you performed using words. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: lot of words, so lets make up the numbers you roll say 6 and 9

OpenStudy (anonymous):



OpenStudy (anonymous):

you want the length of the hypotenuse in english, square 6, square 9, add them together, that gives the square of the hypotenuse, so the hypotenuse is the square root of that number in math \[6^2+9^2=36+81=117\] and so the hypotenuse is \(\sqrt{117}\) whatever that is

OpenStudy (anonymous):

ok now i read that you are only supposed to roll one die, so you cannot roll a 9, lets say you roll at 3 and a 5

OpenStudy (anonymous):



OpenStudy (anonymous):

just like before \[3^2+5^2=9+25=34\] so the hypotenuse is \(\sqrt{34}\)

OpenStudy (anonymous):

ok sorry it took me so long to reply. Thank you for your help.

OpenStudy (anonymous):

What do I do for the second part. The easy part is step one all the heavy lifting is in step 2 & 3 @satellite73